# CUE-RAG: Towards Accurate and Cost-Efficient Graph-Based RAG via Multi-Partite Graph and Query-Driven Iterative Retrieval

### v1

@@ 4.3. Effectiveness of Hybrid Extraction (Q2)
Table 3. Performance comparison of CUE-RAG-0.5 using different core chunk selection strategies. The table highlights the best and second-best results in each column with bold and underlined formatting, respectively.
@@
CUE-RAG-0.5 (RANDOM) 25.18 21.32 54.44 57.60 34.00 46.90 36.61 33.93 61.59 63.40 49.49 54.70 43.55 46.31
CUE-RAG-0.5 (COS) 24.64 21.42 55.80 59.10 35.18 49.10 36.19 32.13 63.65 64.90 52.83 59.10 44.72 47.63
CUE-RAG-0.5 (BLEU) 25.60 22.12 55.80 59.50 35.52 48.90 36.68 32.93 63.44 64.10 52.49 58.80 44.92 47.73
CUE-RAG-1.0 (BLEU) 25.48 22.32 55.97 59.20 37.41 50.70 36.50 32.83 64.02 64.70 55.20 61.00 45.76 48.46
@@
This set of experiments evaluates the effectiveness of our hybrid extraction strategy for knowledge units, as described in Section 3.2. Table 3 reports the QA performance of four methods, each constrained to use only 50% of the total token budget for LLM processing ($\alpha=0.5$). The methods differ in their chunk selection strategies: RANDOM (random selection), COS (our Chunk-Selection using cosine similarity), and BLEU (our Chunk-Selection using BLEU score). For direct comparison, we also include the performance of CUE-RAG-1.0 (BLEU), which uses 100% of the token budget. Results show that both COS and BLEU significantly outperform RANDOM across average F1 and Acc. Among them, the BLEU-based strategy achieves the highest performance, exceeding RANDOM and COS by 3.14% and 0.46% in average F1, and by 3.05% and 0.21% in average Acc, respectively. These gains underscore the effectiveness of our hybrid extraction approach in the offline indexing stage.
@@
While CUE-RAG-1.0 (BLEU) generally yields stronger results, CUE-RAG-0.5 (BLEU) achieves 94.95%–100.51% of its performance, with half the token usage. A particularly interesting phenomenon emerges on the MuSiQue dataset with LLaMA3.0-8B and Qwen2.5-32B, where CUE-RAG-0.5 (BLEU) occasionally outperforms its full-token counterpart. This suggests that certain questions in MuSiQue exhibit a stronger lexical alignment with original corpus sentences than with LLM-extracted knowledge units, as the paraphrasing process may substitute critical terms with less relevant alternatives. Overall, these results demonstrate that our hybrid extraction strategy can maintain robust QA performance while significantly optimizing token efficiency during indexing.
@@ 4.4. Ablation Study (Q3)
Table 4. Ablation study of CUE-RAG-1.0, evaluating the contributions of key components: Knowledge Anchoring (KA.), Spreading Activation (SA.), and Query Updating (QU.). The best performance for each metric is highlighted in bold, while the second-best result is underlined.
@@
CUE-RAG-1.0 25.48 22.32 55.97 59.20 37.41 50.70 36.50 32.83 64.02 64.70 55.20 61.00 45.76 48.46
w/o KA. 25.26 22.02 52.93 55.50 36.75 50.10 32.64 29.13 60.50 62.00 52.58 58.60 43.44 46.23
w/o SA. 23.18 19.52 53.48 55.70 34.76 45.60 31.64 28.03 59.12 59.90 48.57 55.70 41.79 44.08
w/o QU. 24.51 20.72 54.51 56.80 38.00 51.50 35.36 32.03 62.52 64.10 55.27 61.80 45.03 47.83
w/o QU. or KA. 22.53 18.82 52.00 53.90 36.52 49.50 31.93 28.33 60.14 62.00 52.72 58.90 42.64 45.24
w/o QU. or SA. 21.57 17.42 51.22 53.50 33.43 46.10 31.84 27.93 56.71 58.30 48.34 55.80 40.52 43.18
This experiment demonstrates the role of the first two stages, Spreading Activation (SA. below) and Knowledge Anchoring (KA. below) in Section 3.3 and the Query Updating (QU. below) in Iterative Retrieval, as Final Re-reanking is indispensable components of the overall search pipeline. Table 4 shows the performance of removing each mechanism, indicating that both contribute to the improvement of CUE-RAG’s retrieval capability. Specifically, Spreading Activation enhances the average F1 and Acc. by 9.5% and 9.95%, respectively, while Knowledge Anchoring provides improvement of 5.34% and 4.83%, and Query Updating improves the average F1 and Acc. by 1.63% and 1.32%. The superior performance of Spreading Activation can be attributed to its alignment in semantic granularity: the entities extracted from the query by the LLM are matched against entity nodes in the hierarchical graph, ensuring consistency. In contrast, Knowledge Anchoring relies on semantic embeddings of the query to align with knowledge unit nodes in the graph, which may introduce noise due to potential inconsistencies in semantic granularity. Query Updating improves F1 by up to 3.96% and Acc by 7.72% on MuSiQue, and by 2.68% and 4.23% on HotpotQA, but slightly degrades on 2Wiki. The performance decline on 2Wiki stems from its frequent semantically similar entities (e.g., familial relations or homonyms like Elizabeth I/II), which challenge disambiguation. Nevertheless, they collectively enhance the retrieval performance of CUE-RAG.
@@ 4.5. Parameter Sensitivity (Q4)
In the final set of experiments, we analyze the sensitivity of our proposed approach using Qwen2.5-32B with respect to key hyperparameters: the number of top-$K$ retrieved results $K$, beam size $M$, search depth $D$, and token constraint coefficient $\alpha$. We evaluate $K\in\{3,5,7,9\}$, $M\in\{3,5,7,9\}$, $D\in\{1,2,3,4\}$, and $\alpha\in\{0,0.25,0.5,0.75,1.0\}$.
First, as shown in Figure 6 (a), CUE-RAG does not exhibit a consistent improvement in generation quality with increasing $K$ across all three datasets. Instead, the results suggest that $K=3$ provides sufficient retrieval results, while larger values introduce noise without meaningful gains. Next, Figure 6 (b) reveals that generation quality generally improves with larger beam sizes $M$. However, the gains plateau beyond $M=5$ or $M=7$, and a slight decline occurs at $M=9$. This indicates that $M=5$ strikes an optimal balance between generation quality and computational efficiency, as excessively large beam sizes may introduce noise with diminishing returns. Furthermore, Figure 6 (c) demonstrates that deeper search depth $D$ leads to better performance, which aligns with the nature of the datasets, where most questions require 2 hops while fewer necessitate 3-4 hops. Nevertheless, when $D=4$, generation quality slightly deteriorates, likely due to over-retrieval of irrelevant information for most 2-hops questions. Finally, Figure 6 (d) shows a positive correlation between the token constraint coefficient $\alpha$ and generation quality. This is expected because higher $\alpha$ values permit the LLM to process more chunks, thereby improving the quality of knowledge units and enhancing retrieval accuracy.
3333555577779999323232323434343436363636ValueMuSiQue3333555577779999636363636464646465656565(a) Top-K, default 3HotpotQA33335555777799995252525257575757626262622Wiki3333555577779999313131313434343437373737Value3333555577779999606060606464646468686868(b) Beam Size M, default 533335555777799995050505057.557.557.557.5656565651111222233334444202020203030303040404040Value1111222233334444555555556060606065656565(c) Search Depth D, default 31111222233334444505050505656565662626262000.250.250.250.250.50.50.50.50.750.750.750.751111303030303434343438383838Value000.250.250.250.250.50.50.50.50.750.750.750.751111626262626464646466666666(d) Token constraint $\mathbf{\alpha}$, default 1F1Acc.000.250.250.250.250.50.50.50.50.750.750.750.751111494949495555555561616161
Figure 6. Performance analysis under different hyperparameter settings.
@@ 5. Related Works
We review some indexing and retrieval strategies in graph-based RAG methods. For a comprehensive discussion, we refer the reader to previous surveys (Peng et al., 2024; Zhang et al., 2025) and an in-depth analysis of graph-based RAG in a unified framework (Zhou et al., 2025).
@@
Among the early approaches, RAPTOR (Sarthi et al., 2024) introduces a bottom-up tree construction strategy. It begins with raw text chunks as leaf nodes, clusters their embeddings via Gaussian Mixture Models (McInnes et al., 2018), then recursively summarizes clusters using LLMs and re-clusters them to build the tree bottom-up. During retrieval, all tree nodes are indexed in a vector database, enabling a fast semantic similarity search with query embeddings. Building on RAPTOR, SIRERAG (Zhang et al., 2024) introduces a dual tree framework. It extracts propositions and entities from texts using LLMs, regrouping propositions linked to the same entity into new passages. These passages are then organized into a relatedness tree alongside the RAPTOR similarity tree, with both trees jointly indexed in the same vector database for retrieval. This hybrid approach aims to balance semantic similarity and relatedness coherence. Although these tree-based approaches demonstrate effective hierarchical organization, recent works have explored graph-based alternatives for richer structural representation. Microsoft’s GraphRAG (Edge et al., 2024) first constructs a text-attributed KG by extracting entities, relationships, and other detailed contextual features from text fragments through LLMs. KG is then hierarchically clustered with the Leiden algorithm (Traag et al., 2019) , and each cluster is summarized into a community report. For retrieval, the proposed local search retrieves relevant context from entities, relationships, community reports, and text chunks to augment response. LightRAG (Guo et al., 2024) constructs a KG from text chunks while using LLMs to augment both entities and relations with extracted keywords and leverages LLMs to extract query-relevant low-/high-level keywords for retrieval. It integrates local, global, and hybrid search strategies to fetch matching entities and relations, which are then combined with text chunks for answer generation. KETRAG (Huang et al., 2025) selects core text chunks based on their PageRank centrality in a KNN graph built using both semantic and lexical similarity. It then constructs a KG following Microsoft’s GraphRAG, though without community detection or report summarization, while organizing the remaining text chunks into a text-keyword bipartite graph. For retrieval, it adopts standard local search but innovates by extracting ego networks from both entity and keyword channels to improve LLM generation. HippoRAG (Gutiérrez et al., 2024) constructs a KG by extracting triplets from the text chunks using LLMs and enhances its quality by linking similar entities via semantic embeddings. For multi-hop reasoning, it employs PPR to retrieve relevant KG entities and augments the response with entities’ associated text chunks for LLM generation.
@@ 6. Conclusion
In this work, we propose CUE-RAG, a graph-based RAG approach that features a multi-partite graph index integrating chunks, knowledge units, and entities. To efficiently build this index, we introduce a hybrid extraction strategy for knowledge unit that maximizes LLM processing benefits while minimizing token usage. For retrieval, we design a query-driven iterative retrieval (Q-Iter) that ensures relevant retrieval results. These designs enhance graph quality and retrieval relevance while reducing LLM token consumption. Experimental results on three QA benchmarks demonstrate that CUE-RAG significantly outperforms SOTA baselines in both QA performance and cost efficiency. Notably, its zero-token variant achieves comparable or superior performance, highlighting the robustness and effectiveness of our approach.
@@ Appendix A Supplementary Details
Table 5 presents the key statistics for the three benchmark datasets used in our study: MuSiQue, HotpotQA, and 2Wiki, including corpora size, number of questions, and total token count. Table 6 reports the node counts for the CUE-RAG-1.0 graph index, detailing the quantities of text nodes, knowledge unit nodes, and entity nodes for each dataset. Algorithm 4 provides the complete pseudocode for the Entity-Anchoring procedure. Figure 7 and 8 illustrate the LLM instruction prompts used for extracting knowledge units and answer generations, respectively.

### v3

@@ 4.2. Overall performance evaluation (Q1)
Table 3. Performance comparison of Clue-RAG-0.5 using different core chunk selection strategies. The table highlights the best and second-best results in each column with bold and underlined formatting, respectively.
@@
Clue-RAG-0.5 (RANDOM) 25.18 21.32 54.44 57.60 34.00 46.90 36.61 33.93 61.59 63.40 49.49 54.70 43.55 46.31
Clue-RAG-0.5 (COS) 24.64 21.42 55.80 59.10 35.18 49.10 36.19 32.13 63.65 64.90 52.83 59.10 44.72 47.63
Clue-RAG-0.5 (BLEU) 25.60 22.12 55.80 59.50 35.52 48.90 36.68 32.93 63.44 64.10 52.49 58.80 44.92 47.73
Clue-RAG-1.0 (BLEU) 25.48 22.32 55.97 59.20 37.41 50.70 36.50 32.83 64.02 64.70 55.20 61.00 45.76 48.46
In the first set of experiments, we compare Clue-RAG against ten existing solutions in terms of QA performance and cost efficiency. As shown in Table 2, Clue-RAG consistently achieves superior QA performance across multiple benchmark datasets when evaluated with different LLMs, with Clue-RAG-1.0 establishing SOTA performance in both average F1 and Acc. Specifically, it outperforms the strong baseline KETRAG by 21.53% in average F1, while also surpassing another competitive baseline SIRERAG by 5.75% in average Acc. Besides, Clue-RAG-0.5 demonstrates competitive QA performance relative to KETRAG, and achieves improvements of 19.29%/27.19% in average F1/Acc. Notably, Clue-RAG-0.0, which entirely avoids LLM usage in indexing, still matches or exceeds the capabilities of two strong baselines, achieving a 13.01%/52.50% average F1 improvement over KETRAG and SIRERAG, respectively. Beyond outperforming these two competitive baselines, Clue-RAG demonstrates superior QA performance across all evaluated datasets compared to the other eight alternatives. This consistent advantage highlights the effectiveness of its innovative indexing and retrieval strategies in multi-hop QA scenery.
Meanwhile, compared to KETRAG, Clue-RAG-0.5 reduces token costs by up to 9.41%/83.87% during offline indexing and online retrieval, as shown in Figure 5 and 6. Remarkably, during offline indexing, Clue-RAG-0.0 reduces token costs by 100% compared to KETRAG and SIRERAG. During online retrieval, it reduces up to 82.92%/56.06% token costs compared to these two strong baselines. Hence, Clue-RAG is token-cost efficient.
[…]
Figure 5. Token cost in offline indexing using LLaMA3.0-8B (Clue-RAG-0.0 requires no LLM).
4.3. Effectiveness of hybrid extraction (Q2)
@@
In this experiment, we evaluate the effectiveness of our hybrid extraction strategy for knowledge units, as described in Section 3.2. Table 3 reports the QA performance of four methods, each constrained to use only 50% of the total token budget for LLM processing ($\alpha=0.5$). The methods differ in their chunk selection strategies: RANDOM (random selection), COS (our Chunk-Selection using cosine similarity), and BLEU (our Chunk-Selection using BLEU score). For direct comparison, we also include the performance of Clue-RAG-1.0 (BLEU), which uses 100% of the token budget. The results show that both COS and BLEU significantly outperform RANDOM on average F1 and Acc. Particularly, the BLEU-based strategy achieves the highest performance, exceeding RANDOM and COS by 3.14% and 0.46% in average F1, and by 3.05% and 0.21% in average Acc, respectively. These gains underscore the effectiveness of our hybrid extraction approach in the offline indexing stage.
@@
While Clue-RAG-1.0 (BLEU) yields stronger results, Clue-RAG-0.5 (BLEU) achieves 94.95%–100.51% of its performance, with half of the token usage. An interesting phenomenon emerges on the MuSiQue dataset with LLaMA3.0-8B and Qwen2.5-32B, where Clue-RAG-0.5 (BLEU) occasionally outperforms its full-token counterpart. This suggests that certain questions in MuSiQue exhibit a stronger lexical alignment with original sentences than with LLM-extracted knowledge units, as the paraphrasing process may substitute critical terms with less relevant alternatives. Overall, these results demonstrate that hybrid extraction strategy can maintain robust QA performance while significantly reducing the token cost.
[…]
Figure 6. Token cost in online retrieval using LLaMA3.0-8B.
@@ 4.4. Ablation study (Q3)
Table 4. Ablation study of Clue-RAG-1.0, evaluating the contributions of three key components (K-A, S-A, and Q-U). The best performance for each metric is highlighted in bold, while the second-best result is underlined.
@@
Clue-RAG-1.0 25.48 22.32 55.97 59.20 37.41 50.70 36.50 32.83 64.02 64.70 55.20 61.00 45.76 48.46
w/o K-A 25.26 22.02 52.93 55.50 36.75 50.10 32.64 29.13 60.50 62.00 52.58 58.60 43.44 46.23
w/o S-A 23.18 19.52 53.48 55.70 34.76 45.60 31.64 28.03 59.12 59.90 48.57 55.70 41.79 44.08
w/o Q-U 24.51 20.72 54.51 56.80 38.00 51.50 35.36 32.03 62.52 64.10 55.27 61.80 45.03 47.83
w/o Q-U or K-A 22.53 18.82 52.00 53.90 36.52 49.50 31.93 28.33 60.14 62.00 52.72 58.90 42.64 45.24
w/o Q-U or S-A 21.57 17.42 51.22 53.50 33.43 46.10 31.84 27.93 56.71 58.30 48.34 55.80 40.52 43.18
In this experiment, we evaluate the key components of Q-Iter, i.e., Spreading Activation (S-A below) and Knowledge Anchoring (K-A below) of Step-1, and the Query Updating (Q-U below) of Step-2. Table 4 shows the performance of removing each component. We observe that S-A enhances the average F1 and Acc. by 9.5% and 9.95%, respectively, while K-A provides improvement of 5.34% and 4.83%, and Q-U improves the average F1 and Acc. by 1.63% and 1.32%. The superior performance of S-A can be attributed to its alignment in semantic granularity: The entities extracted from the query by the LLM are matched against entity nodes in the multi-partite graph, ensuring consistency. In contrast, K-A relies on semantic embeddings of the query to align with knowledge unit nodes in the graph, which may introduce noise due to potential inconsistencies in semantic granularity. Q-U improves F1 by up to 3.96% and Acc by 7.72% on MuSiQue, and by 2.68% and 4.23% on HotpotQA, but slightly degrades on 2Wiki. The performance decline on 2Wiki stems from its frequent semantically similar entities (e.g., family relations or homonyms like Elizabeth I/II), which challenge disambiguation. Nevertheless, they collectively enhance the retrieval performance of Clue-RAG.
@@ 4.5. Parameter sensitivity (Q4)
In this experiment, we analyze the sensitivity of Clue-RAG using Qwen2.5-32B w.r.t. key hyperparameters: the number $K$ of top-$K$ retrieved results, beam size $M$, search depth $D$, and token constraint coefficient $\alpha$, where $K\in\{3,5,7,9\}$, $M\in\{3,5,7,9\}$, $D\in\{1,2,3,4\}$, and $\alpha\in\{0,0.25,0.5,0.75,1.0\}$.
First, as shown in Figure 7 (a), Clue-RAG does not exhibit a consistent improvement in generation quality with increasing $K$ across all three datasets. Instead, the results suggest that $K=3$ provides sufficient retrieval results, while larger values introduce noise without meaningful gains. Next, Figure 7 (b) reveals that generation quality generally improves with larger beam sizes $M$. However, the gains plateau beyond $M=5$, and a slight decline occurs at $M=7,9$. This indicates that $M=5$ strikes an optimal balance between generation quality and computational efficiency, as excessively large beam sizes may introduce noise with diminishing returns. Furthermore, Figure 7 (c) demonstrates that deeper search depth $D$ leads to better performance, which aligns with the nature of the datasets, where most questions require 2 hops while fewer necessitate 3-4 hops. Nevertheless, when $D=4$, generation quality slightly deteriorates, likely due to over-retrieval of irrelevant information for most 2-hops questions. Finally, Figure 7 (d) shows a positive correlation between the token constraint coefficient $\alpha$ and generation quality. This is expected because higher $\alpha$ values allow the LLM to process more chunks, thereby improving the quality of knowledge units and enhancing retrieval accuracy. Consistent results are observed across other datasets in Appendix Figure 8.
33557799313134343737(a) Top-K33557799313134343737(b) M11223344202030304040(c) DMuSiQueF1Acc.00.250.250.50.50.750.7511303034343838(d) $\mathbf{\alpha}$
Figure 7. Performance analysis under different hyperparameter settings on MuSiQue dataset.
@@ 5. Related Works
In this section, we review the representative RAG methods based on vector database (VDB) (Pan et al., 2024; Zhang et al., 2024b), trees, and graphs. For more details, please see recent surveys (Peng et al., 2024; Zhang et al., 2025a; Fan et al., 2024) and empirical study (Zhou et al., 2025).
$\bullet$ VDB-based RAG. Vallina RAG is a basic RAG solution that (1) splits the corpora into text chunks, (2) encodes them into embeddings via an embedding model, and (3) stores embeddings in a VDB. During online retrieval, the same model embeds the query and then retrieves top-$K$ relevant text chunks for augmented generation.
$\bullet$ Tree-based RAG. RAPTOR (Sarthi et al., 2024) introduces a tree-based index, which can be built in a bottom-up tree manner. Specifically, it begins with raw text chunks as leaf nodes, clusters their embeddings via Gaussian Mixture Models (McInnes et al., 2018), then recursively summarizes clusters using LLMs and re-clusters them to build the tree bottom-up. During online retrieval, all tree nodes are indexed in a VDB, enabling a fast semantic similarity search with query embeddings. Building on RAPTOR, SIRERAG (Zhang et al., 2024a) introduces a dual tree framework, which first extracts propositions and entities from texts using LLMs, regrouping propositions linked to the same entity into new passages, and then organize these passages into a relatedness tree alongside the RAPTOR similarity tree, with both trees jointly indexed in the same vector database for retrieval. This hybrid approach aims to balance semantic similarity and relatedness coherence. EraRAG (Zhang et al., 2025b) is a novel hierarchical tree-based RAG method that focuses on the dynamic scenario, allowing for efficient and incremental index updates as the corpus evolves.
@@
$\bullet$ Graph-based RAG. Compared to trees, graphs are more effective for modeling complicated relationships. Microsoft’s GraphRAG (Edge et al., 2024) first constructs a text-attributed KG by extracting entities, relationships, and other detailed contextual features from text fragments through LLMs. Then, it employs Leiden algorithm (Traag et al., 2019) to cluster the KG into some hierarchical clusters, where each cluster is associated with a community report. For retrieval, it uses the local search to retrieve relevant context from entities, relationships, community reports, and text chunks to augment response. ArchRAG (Wang et al., 2025) organizes attributed communities hierarchically, and augments the question using summaries of attributed communities. LightRAG (Guo et al., 2024) constructs a KG from text chunks while using LLMs to augment both entities and relations with extracted keywords and leverages LLMs to extract query-relevant low-/high-level keywords for retrieval. KETRAG (Huang et al., 2025) first selects core text chunks from a KNN graph built using both semantic and lexical similarity, and then constructs a KG. It supports online retrieval by using the standard local search but extracting ego networks to improve LLM generation. HippoRAG (Gutiérrez et al., 2024) constructs a KG by extracting triplets from the text chunks using LLMs and enhances its quality by linking similar entities via semantic embeddings. For multi-hop reasoning, it employs PPR to retrieve relevant KG entities and augments the response with entities’ associated text chunks for LLM generation.
@@ 6. Conclusion
In this work, we propose Clue-RAG, a graph-based RAG approach that features a multi-partite graph index integrating chunks, knowledge units, and entities. To efficiently build this index, we introduce a hybrid extraction strategy for knowledge unit that maximizes LLM processing benefits while minimizing token usage. For online retrieval, we design a query-driven iterative retrieval (Q-Iter) that ensures relevant retrieval results. These designs enhance graph quality and retrieval relevance while reducing LLM token consumption. Experimental results on three QA benchmarks demonstrate that Clue-RAG significantly outperforms SOTA baselines in both QA performance and cost efficiency. Notably, its zero-token variant achieves comparable or superior performance, highlighting the robustness and effectiveness of our approach. In the future, we will extend our Clue-RAG for processing multimodal data.
@@ Appendix A Supplementary Details
A.1. Experimental Settings
For token cost comparison between offline indexing and online retrieval using LLaMA3.0-8B shown in Figure 5 and 6, we account for both prompt and completion tokens. For methods requiring LLM-based query preprocessing (e.g., keyword or entity extraction), these additional token costs are included in our calculations. During the LLM generation phase, we employ a strategy by duplicating the input query to the end of the original prompt. This design explicitly mitigates the ”lost in the middle” effect (Liu et al., 2023) observed in approaches like LightRAG, where placing the question before lengthy context can cause the model to overlook the task objective (Ram et al., 2023; Liu et al., 2023). Furthermore, inspired by the prompt designs of SIRERAG (Zhang et al., 2024a) and KETRAG (Huang et al., 2025), which include phrases like ”Answer this question as short as possible” to prevent verbose or irrelevant outputs, we uniformly add this instruction to the final prompt of each method. This ensures concise responses, maintaining evaluation fairness and highlighting our method’s effectiveness. For other hyperparameters of each method, we follow the original settings in their available codes. For NLP pre-processing, we employ the following tools: cl100k_base from tiktoken library (OpenAI, [n. d.]) for word tokenization and counting, NLTK’s sent_tokenize (Bird et al., 2009) for sentence segmentation, and spaCy (Honnibal et al., 2020) for named entity recognition. For LLM inference, we use LLaMA3.0-8B (Grattafiori et al., 2024) and Qwen2.5-32B (Bai et al., 2023), with a maximum token limit of 8,000. In top-$K$ selection tasks, we set $K=5$ to accommodate token constraints. The embedding model for vector search is BGE-M3 (Chen et al., 2024), while re-ranking is performed using the lightweight yet powerful BGE-Reranker-v2-M3 (Chen et al., 2024). As for our proposed approach, Clue-RAG employs the following default configurations: the number of most relevant results $K$ is set to 3, search depth $D$ to 3, beam size $M$ per depth to 5, and returned candidate chunks $N$ to 5, ensuring that all methods ultimately retrieve the same number of query-relevant chunks (i.e., 5) for fair comparison. By default, we set the value of $\alpha$ to 1 if it is not explicitly specified. Additionally, BLEU score (Papineni et al., 2002) serves as the core selection metric for relevance scoring.
A.2. Dataset Details
Table 5 presents the key statistics for the three benchmark datasets used in our study: MuSiQue, HotpotQA, and 2Wiki, including corpora size, number of questions, and total token count. Table 6 reports the node counts for the Clue-RAG-1.0 graph index, detailing the quantities of text nodes, knowledge unit nodes, and entity nodes for each dataset. Figure 8 presents the performance under different hyperparameter settings on HotpotQA and 2Wiki datasets.